# KG-Infused RAG: Augmenting Corpus-Based RAG with External Knowledge Graphs

### v1

@@ F.1 Experimental Result of Self-Ask
As shown in Table 21, combining KG-Infused RAG with Self-Ask yields moderate gains on four datasets, while a slight performance drop is observed on HotpotQA. Compared to Self-RAG, the overall improvements are smaller and less consistent. This difference is likely due to Self-Ask’s step-by-step decomposition of complex questions into simpler single-hop sub-questions (see prompt in Table 18 and examples in Table 24). This decomposition enables effective retrieval using only the corpus, thereby reducing reliance on external structured knowledge. In such cases, the marginal benefit of KG-based augmentation diminishes, and injecting external facts may introduce irrelevant information that interferes with reasoning.
@@
Appendix G Limitation
KG Data Modalities.
This work focuses on activating structured knowledge in the form of triples from KGs, which we integrate into a corpus-based RAG pipeline. However, KGs often contain heterogeneous forms of information beyond triples, such as long-form text, images, and tables. Our current framework is designed specifically for triple-based activation and does not yet account for these heterogeneous modalities. Effectively leveraging such diverse KG content in the activation process is an important direction for future research, potentially enabling broader applicability and richer factual grounding.
Triple Utilization.
Our method leverages KG triples in two main ways: query expansion and knowledge augmentation of retrieved passages. While effective, we believe KG facts can be further exploited through other strategies to better align with corpus retrieval and enhance downstream reasoning. Investigating more flexible or task-specific integration mechanisms remains an open and valuable area of future work.
@@ Appendix H Case Study
H.1 Case I of Main Results: Direct Answer Retrieval via KG Activation
@@
Table 22 presents a case study comparing the answer generation process of KG-Infused RAG and two baselines. This example illustrates how, through two rounds of KG-guided activation, KG-Infused RAGdirectly acquires answer-relevant triples from the KG, providing sufficient structured context for the subsequent stages.
@@
In contrast, the answer generation processes of the two baselines suffer from the lack of such structured knowledge input. Vanilla RAG mistakenly associates the missile “AAM-N-4 Oriole” with the unrelated “Oriole Records” label due to misleading surface-form overlaps in the corpus, leading to an erroneous inference that CBS followed the missile manufacturer. Meanwhile, Vanilla QE misinterprets the earlier development date of Ryan Aeronautical’s missile project as indicating that Ryan "followed" the Glenn L. Martin Company, incorrectly equating temporal precedence with corporate succession.
@@
H.2 Case II of Main Results: KG-Guided Query Simplification via Intermediate Nodes
@@
Even when the answer node itself is not present in the KG—making it impossible to directly retrieve the answer via multi-round KG-guided activation—the retrieved multi-hop intermediate nodes can still play a crucial role. Specifically, KG-guided Spreading Activation can still identify intermediate entities and relations that form partial reasoning chains, providing useful cues for query reformulation and retrieval. In such cases, KG-Based QE transforms the original complex question into one or more simplified single-hop sub-questions that focus on intermediate concepts surfaced by the KG. These simplified queries improve corpus retrieval by aligning better with the surface forms and local context found in text passages. Table 23 presents several examples of simplified single-hop questions generated by KG-Based QE, demonstrating how intermediate KG nodes help bridge the gap between the original query and corpus evidence.
H.3 Case of Self-Ask

### v2

@@ F.1 Results with Self-Ask
As shown in Table 23, combining KG-Infused RAGwith Self-Ask yields moderate gains on four datasets, but a slight performance drop is observed on HotpotQA. Compared with Self-RAG, the overall improvements are smaller and less consistent. This difference is likely due to Self-Ask’s step-by-step decomposition of complex questions into simpler single-hop sub-questions (see the prompt in Table 19 and examples in Table 26). This decomposition enables effective corpus-only retrieval, thereby reducing reliance on external structured knowledge. In such cases, the marginal benefit of KG-based augmentation diminishes, and injecting external facts may introduce irrelevant information that interferes with reasoning.
Table 22: Cost of KG construction and inference (Total tokens = Input + Output).
KG Construction
Method Input tokens Output tokens Total tokens Time
GraphRAG 11,853s
LightRAG 141,900,133 26,337,116 168,237,249 31,905s
KG2RAG 4,029,418 985,660 5,015,078 6,105s
KG-Infused RAG / / / /
Inference
Method Input tokens Output tokens Total tokens Time
GraphRAG (local) 4,469.97 9.56 4,479.53 5.38s
GraphRAG (global) 282,222.93 2,449.92 284,672.85 32.69s
LightRAG (local) 17,397.29 71.28 17,468.57 4.41s
LightRAG (global) 12,281.29 80.09 12,361.38 3.49s
KG2RAG 4.31s
KG-Infused RAG (GPT-4o-mini) 3,543.06 560.35 4,103.41 3.98s
F.2 Computation Cost Analysis
Table 22 reports the token usage and runtime of different KG-based methods during both KG construction and inference. For GraphRAG and LightRAG, we use the default gpt-4o-mini model for knowledge-graph construction as well as inference. To quantify inference overhead, we report the average cost of each method on the MuSiQue dataset.
@@ F.2 Computation Cost Analysis
Appendix G Limitations
Heterogeneous KG Modalities.
This work focuses on activating structured knowledge from knowledge graphs (KGs) in the form of triples, and we integrate these into a corpus-based retrieval-augmented generation (RAG) pipeline. However, KGs often contain heterogeneous information beyond triples, such as long-form text, images, and tables. Our current framework is designed specifically for triple-based activation and does not yet handle these heterogeneous modalities. Effectively leveraging such diverse KG content in the activation process is an important direction for future research, potentially enabling broader applicability and richer factual grounding.
KG Triple Utilization.
Our method leverages KG triples in two main ways: query expansion and knowledge augmentation for retrieved passages. While effective, KG facts could be further exploited via additional strategies that better align with corpus retrieval and enhance downstream reasoning. Investigating more flexible or task-specific integration mechanisms remains an open and valuable area of future work.
@@ Appendix H Case Study
H.1 Case I: Direct Answer Retrieval via KG-Guided Spreading Activation
@@
Table 24 presents a case study comparing the answer-generation process of KG-Infused RAG with two baselines. This example illustrates how, through two rounds of KG-guided spreading activation, KG-Infused RAG directly acquires triples relevant to the answer from the KG, providing sufficient structured context for subsequent stages.
@@
In contrast, the answer-generation processes of the two baselines suffer from a lack of such structured input. Vanilla RAG mistakenly associates the missile “AAM-N-4 Oriole” with the unrelated “Oriole Records” label due to misleading surface-form overlaps in the corpus, leading to an erroneous inference that CBS followed the missile’s manufacturer. Meanwhile, Vanilla QE misinterprets the earlier development date of Ryan Aeronautical’s missile project as indicating that Ryan “followed” the Glenn L. Martin Company, incorrectly equating temporal precedence with corporate succession.
@@
H.2 Case II: KG-Guided Query Simplification via Intermediate Nodes
@@
Even when the answer node is not present in the KG—making it impossible to directly retrieve the answer via multi-round KG-guided activation—multi-hop intermediate nodes can still play a crucial role. Specifically, KG-Guided Spreading Activation can identify intermediate entities and relations that form partial reasoning chains, providing useful cues for query reformulation and retrieval.
In such cases, KG-Based QE transforms the original complex question into one or more simplified single-hop sub-questions that focus on intermediate concepts surfaced by the KG. These simplified queries improve corpus retrieval by aligning better with the surface forms and local context found in text passages. Table 25 presents several examples of simplified single-hop questions generated by KG-Based QE, demonstrating how intermediate KG nodes help bridge the gap between the original query and corpus evidence.
H.3 Case III: Self-Ask with KG-Infused RAGas a Plug-in